A sales manager for an advertising agency believes there is a relationship between the number of contacts and the amount of the sales. To verify this, the following data was collected:
Sales ($)   Contacts
25,000   15
16,000   15
28,000   20
60,000   18
80,000   46
35,000   24
85,000   45
80,000   45
95,000   55
110,000   52
Which is the:
Independent variable?   ___________________

Dependent variable?   ___________________

Calculate the coefficient of correlation and interpret the results.

Determine both the coefficient of determination and non-determination and explain the result.


Determine the unrounded regression equation for this data set.

Using the regression equation predict the amount of sales when 50 contacts are made. Show all work involved in finding your answer.
Answer #1

Here independent variable is Number of contacts

Dependent variable is Sales (In thousands)

Here coefficient of correlation is calculated with the help of values of , , , ,

Contacts(x) Sales ($)(y) x^2 xy y^2
15 25 225 375 625
15 16 225 240 256
20 28 400 560 784
18 60 324 1080 3600
46 80 2116 3680 6400
24 35 576 840 1225
45 85 2025 3825 7225
45 80 2025 3600 6400
55 95 3025 5225 9025
52 110 2704 5720 12100
Sum 335 614 13645 25145 47640

Correaltion coefficient r = [n(Σxy) - (Σx)((Σy)] / sqrt [ (n (Σx2 ) - (Σx)2) ( n (Σy2 ) - (Σy)2)]

r = [10 * 25145 - 335 * 614]/ [(10 * 13645 - 335 * 335) * (10 * 47640 - 335 * 335)]

r = 0.9325

Here as the correlation coefficient is near to 1 that means there is strong positive correlation between number of contacts and sales.

Here let say the regression equation is

y^ = a + bx

a = [(Σy) (Σx2 ) - (Σx) (Σxy)]/ [ n (Σx2 ) - (Σx)2 ]

a = [614 * 13645 - 335 * 25145]/ [10 * 13645- 335 * 335]

a = -1.88

b = [ n(Σxy) - (Σx)((Σy)]/ [ n (Σx2 ) - (Σx)2 ]

b = [10 * 25145 - 335 * 614]/ [10 * 13645- 335 * 335]

b = 1.889

y^ = -1.88 + 1.889x

where x = number of contacts

y^ = predicted value of sales (in thousands dollars)

Here if x = 50

y^ = -1.88 + 1.889 * 50

y^ = $92568 is  the amount of sales when 50 contacts are made.